OpenStudy (katecc379):

What are the explicit equation and domain for an arithmetic sequence with a first term of 5 and a second term of 2?

OpenStudy (katecc379):

@confluxepic

OpenStudy (welshfella):

common difference d = 2-5 = -3

OpenStudy (welshfella):

an = a1 + (n-1)d

OpenStudy (katecc379):

I've tried that but it isn't any of the options none have the different at the end

OpenStudy (katecc379):

would ti be 5-3(n-1) ?

OpenStudy (welshfella):

it doesn't matter if d is after the parentheses or before an = a1 = d(n-1) is the same

OpenStudy (welshfella):

yes correct

OpenStudy (welshfella):

* a1 +

OpenStudy (welshfella):

OK?

OpenStudy (katecc379):

so then the domain would be all greater than or equal to 1 correct?

OpenStudy (welshfella):

n= 1, 2,3.... Yes

OpenStudy (welshfella):

n>= 1 or in interval notation its [1, +Infinity)

OpenStudy (katecc379):

can you help with another?